# ToonifyGB: StyleGAN-based Gaussian Blendshapes for 3D Stylized Head Avatars

### v2

@@ Introduction
Dataset justin malte_1 nf_01 bala wojtek_1 person_0004
@@
Table 3: Quantitative comparison of 3D head avatars: We evaluate our method and SOTA methods on the INSTA (Zielonka, Bolkart, and Thies 2023) dataset. In each metric group, the best value is highlighted in bold, and the second‑best is underlined.
@@
[…]
INSTA 31.66 0.9591 27.44 0.9159 26.45 0.8937 29.53 0.8896 31.36 0.9452 25.44 0.8478
PointAvatar 30.40 0.9373 24.98 0.8853 25.25 0.8919 27.88 0.8658 28.82 0.9192 23.29 0.8576
FLARE 29.10 0.9363 25.93 0.8973 25.97 0.9027 27.20 0.8761 27.84 0.9216 25.53 0.9015
SplattingAvatar 30.93 0.9482 27.66 0.9243 27.08 0.9202 32.14 0.9272 29.54 0.9400 26.49 0.9075
FlashAvatar 32.16 0.9611 27.45 0.9326 28.02 0.9326 30.27 0.8494 32.02 0.9509 25.49 0.8996
3DGB 32.63 0.9643 28.65 0.9432 28.06 0.9340 33.29 0.9457 32.57 0.9623 23.66 0.8449
@@
Method Metric justin malte_1 nf_01 bala wojtek_1 person_0004
@@
Quantitative Comparison
Video Stabilization
@@
To evaluate the impact of preprocessing, we apply a standard face alignment technique based on a facial keypoint predictor (Kazemi and Sullivan 2014) to the input videos. We compare the original input videos (Original Input, OI) with their aligned counterparts (Aligned Input, AI). Likewise, we compare the “Arcane” style outputs generated from unaligned inputs (Ours Arcane, OA) with those generated from aligned inputs (Aligned Arcane, AA). As shown in Table 2, both the Inter-frame Transformation Fidelity (ITF) and Inter-frame Similarity Index (ISI) scores for AI are consistently lower than those for OI. Similarly, AA exhibits lower ITF and ISI scores compared to OA. These results suggest that applying face alignment and cropping prior to frame-by-frame generation (i.e., AI and AA) tends to introduce greater temporal instability, resulting in more jittery outputs.
3D Head Avatar
We evaluate our method and state‑of‑the‑art methods using standard metrics for animatable head reconstruction, as well as training and rendering times. The quantitative results are presented in Table 3, and the training and rendering times are recorded in Table 4. With the additional integration of stylization, our method achieves performance comparable to the state‑of‑the‑art on the PSNR and SSIM metrics in most cases, and even outperforms them on certain data. Specifically, our method outperforms all other methods on synthesizing the “Arcane” style for the “bala” and “person_0004” data, as well as the “Pixar” style for the “justin” and “nf_01” data.
In addition, although our method integrates stylization into 3D head avatars, its training and rendering times remain comparable to those of the method of 3DGB (Ma et al. 2024). In certain cases, our method is even more efficient in both training and rendering. Combined with the additional time required for video generation (as shown in Table 1), the overall time cost of our method remains acceptable.
@@
Figure 5: Qualitative comparison of baseline and ours: We present 3D head avatars using Gaussian blendshapes synthesized by 3DGB (Ma et al. 2024) and our method.
Qualitative Comparison
We present the original video head frames, the corresponding stylized video frames generated by our method, and the 3D stylized head avatars synthesized using Gaussian blendshapes. The qualitative comparison is shown in Figure 4. The examples are selected from the “bala” dataset in the “Pixar” style and the “wojtek_1” dataset in the “Arcane” style. In the stylized video, the “bala” data exhibits artifacts along the side edge of the head. We attribute this to the latent space distribution learned by StyleGAN, which tends to produce striped artifacts when the viewing angle falls outside the distribution covered by the training data. Notably, these artifacts are not present in the corresponding 3D stylized head avatars rendered by our method. Furthermore, the 3D stylized head avatars successfully preserve fine details from the stylized videos, such as the mole near the eye in the “wojtek_1” dataset. However, since the 3D avatar synthesis mainly focuses on the facial region, the neck area is typically blurred, as observed in both cases. This blurring leads to the lower quantitative performance, since the neck region is included in the evaluation.
The qualitative comparison with 3DGB (Ma et al. 2024) is presented in Figures 5. Our method effectively captures and preserves high-frequency details in the stylized videos. Compared to the SOTA method, ToonifyGB can synthesize 3D stylized head avatars with comparable quality and detail.
@@
Visualization
Generated Video Details
To better demonstrate the visual quality of our generated videos, we present several examples in Figure 2, and select two representative videos for detailed comparison in Figure 3. Specifically, we show real head frames from the “bala” and “wojtek_1” videos, as well as the corresponding heads of generated videos in the “Arcane” style. The results demonstrate that key facial features, such as the beard, mouth shape, and even small details like the black mole above the eye in the lower right image, are well preserved after the stylization process. These details highlight the excellent performance of our method in terms of detail preservation and identity consistency.
Ablation Study
Face Alignment and Cropping
@@
Source Videos for Driving Animation
Compared with the architecture of 3DGB (Ma et al. 2024) that synthesizes 3D photo-realistic head avatars, our framework includes an additional Stage 1 to generate the stylized video. To demonstrate the importance of the generated stylized video in driving the animation, we compare the results of using the original input video (real face) versus our generated stylized video as the driving source, as shown in Figure 6.
@@
Figure 7: Limitation: We present side-view renderings synthesized by 3DGB (Ma et al. 2024) and our method.
Limitation
Our method struggles to render side views of 3D stylized head avatars when the training data (i.e., input video) lacks side-view representations of the real head. As shown in Figure 7, we present side-view renderings synthesized by both 3DGB (Ma et al. 2024) and our method, and this limitation is also observed in the state-of-the-art methods. In fact, existing NeRF-based and Gaussian-based methods have yet to effectively address this issue. Rendering novel views from single-view training data remains an open problem for future research. Two directions to address this limitation include employing 2D GANs to synthesize videos with side views as additional training data, and enhancing the generalization ability of our model.
Conclusion
We propose a novel two-stage framework, ToonifyGB, which utilizes Gaussian blendshapes to synthesize head animations in diverse styles from monocular videos. In Stage 1, the proposed method adopts an improved StyleGAN-based model to generate stylized videos without requiring face alignment or cropping as preprocessing. This results in more temporally stable outputs, providing a reliable foundation for high-quality 3D head avatar animation synthesis. Stage 2 focuses on constructing 3D stylized head avatars using Gaussian blendshapes, enabling fine-grained expression modeling and satisfactory animation. Our method supports real-time generation of stylized avatar animations in popular styles such as “Arcane” and “Pixar”.
For future work, we plan to integrate motion capture technologies to enable real-time expression control of 3D stylized avatars. This direction is expected to further broaden the applicability of ToonifyGB in virtual character interaction and personalized avatar generation.

### v3

@@ 4.1 Baselines
Video Samples justin malte_1 nf_01 bala wojtek_1 person_0004
@@
4.2 Dataset
We evaluate both our method and state-of-the-art photo-realistic avatar synthesis methods using six videos from the INSTA [60] dataset. Each video is cropped and resized to $512\times 512$ resolution, with sequence lengths ranging from 1,000 to 4,000 frames. Following the method of 3DGB [29], we retain the final 350 frames of each video for testing. Both 3DGB [29] and our method apply the same preprocessing pipeline [59, 60], including background removal and FLAME parameter extraction.
Table 3: Quantitative comparison of 3D head avatars: We evaluate our method and state-of-the-art methods on the INSTA [60] dataset. In each metric group, the best value is highlighted in bold, and the second‑best is underlined.
@@ 4.2 Dataset
$uparrow PSNR\uparrow SSIM\uparrow$ $uparrow PSNR\uparrow SSIM\uparrow$ $uparrow PSNR\uparrow SSIM\uparrow$
INSTA [60] 31.66 0.9591 27.44 0.9159 26.45 0.8937 29.53 0.8896 31.36 0.9452 25.44 0.8478
PointAvatar [58] 30.40 0.9373 24.98 0.8853 25.25 0.8919 27.88 0.8658 28.82 0.9192 23.29 0.8576
FLARE [6] 29.10 0.9363 25.93 0.8973 25.97 0.9027 27.20 0.8761 27.84 0.9216 25.53 0.9015
SplattingAvatar [39] 30.93 0.9482 27.66 0.9243 27.08 0.9202 32.14 0.9272 29.54 0.9400 26.49 0.9075
FlashAvatar [45] 32.16 0.9611 27.45 0.9326 28.02 0.9326 30.27 0.8494 32.02 0.9509 25.49 0.8996
3DGB [29] 32.63 0.9643 28.65 0.9432 28.06 0.9340 33.29 0.9457 32.57 0.9623 23.66 0.8449
@@
Video Samples justin malte_1 nf_01 bala wojtek_1 person_0004
@@
4.3 Evaluation Metrics
We employ two metrics to evaluate video stabilization: Inter-frame Transformation Fidelity (ITF) [32, 30, 46] and Inter-frame Similarity Index (ISI) [12, 14]. ITF measures the inter-frame Peak Signal-to-Noise Ratio (PSNR) in dB based on the mean squared error. The intuitive idea of ITF is that a more stable video (i.e., less jittery) will have greater similarity between adjacent frames compared to an unstable version of the same video. ISI computes the average Structural Similarity (SSIM) between adjacent frames across the video. Higher ISI values indicate greater perceptual similarity between frames, leading to improved visual comfort for viewers.
For 3D head avatar synthesis, we evaluate the performance of our method and state-of-the-art methods for photo-realistic avatar synthesis using standard evaluation metrics [56, 29], including Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM). In addition, we record the training time (in minutes) and the rendering speed (in frames per second, fps) for each method. In the ablation study, we additionally adopt the Learned Perceptual Image Patch Similarity (LPIPS) metric to better capture perceptual differences between the synthesized avatars and the ground truth.
4.4 Implementation Details
To ensure a fair performance comparison, the training and testing of all methods are performed on a single RTX 4090 GPU. Our methods are implemented in Python using the PyTorch framework.
For 2D stylized video generation, we use the pre-trained models provided by StyleGANEX [51]. For training the 3D stylized head avatars, we employ the Adam optimizer [23], setting the initial learning rates of the Gaussian properties $\{\mathbf{x}_{k},\alpha_{k},\mathbf{s}_{k},\mathbf{q}_{k},S_{H_{k}}\}$ to $3.2\times 10^{-7}$, $5\times 10^{-5}$, $5\times 10^{-4}$, $1\times 10^{-4}$, and $1.25\times 10^{-3}$, respectively. Following 3DGB [29], the initial number of sampled Gaussians is 50k for the neutral head model and 14k for the mouth interior.
4.5 Quantitative Comparison
4.5.1 Video Stabilization
@@
To evaluate the impact of preprocessing, we apply a standard face alignment technique based on a facial keypoint predictor [21] to the input videos. We compare the original input videos (Original Input, OI) with their aligned counterparts (Aligned Input, AI). Likewise, we compare the “Arcane” style outputs generated from unaligned inputs (Ours Arcane, OA) with those generated from aligned inputs (Aligned Arcane, AA).
As shown in Table 2, both the Inter-frame Transformation Fidelity (ITF) and Inter-frame Similarity Index (ISI) scores for AI are consistently lower than those for OI. Similarly, AA exhibits lower ITF and ISI scores compared to OA. These results suggest that applying face alignment and cropping prior to frame-by-frame generation (i.e., AI and AA) tends to introduce greater temporal instability, resulting in more jittery outputs.
@@
Figure 6: Qualitative comparison of baseline and ours: We present 3D head avatars using Gaussian blendshapes synthesized by 3DGB [29] and our method.
Table 5: User preference study: We conduct user preference studies on the “Arcane” and “Pixar” styles, where users rate their preferences on a scale from 1 to 5, with higher scores indicating greater satisfaction, across three evaluation criteria: Style Consistency, Identity Preservation, and Overall Quality. The highest percentage is highlighted in bold.
Style Arcane Pixar
Evaluation Criteria Style Consistency Identity Preservation Overall Quality Style Consistency Identity Preservation Overall Quality
1: Very Dissatisfied 0.0% 2.8% 0.6% 2.8% 5.6% 2.2%
2: Dissatisfied 5.0% 10.0% 6.1% 5.0% 6.7% 5.0%
3: Neutral 11.1% 15.6% 15.0% 13.3% 24.4% 20.0%
4: Satisfied 32.2% 39.4% 36.7% 29.4% 36.7% 37.8%
5: Very Satisfied 51.7% 32.2% 41.7% 49.4% 26.7% 35.0%
4.5.2 3D Head Avatar
We evaluate our method and state‑of‑the‑art methods using standard metrics for animatable head reconstruction. The quantitative results are presented in Table 3, and the training and rendering times for both the baseline methods and ours are reported in Table 4. With the additional integration of stylization, our method achieves performance comparable to the state‑of‑the‑art on the PSNR and SSIM metrics in most cases, and even outperforms them on certain data. Specifically, our method outperforms all other methods on synthesizing the “Arcane” style for the “bala” and “person_0004” data, as well as the “Pixar” style for the “justin” and “nf_01” data.
In addition, although our method integrates stylization into 3D head avatars, its training and rendering times remain comparable to those of the method of 3DGB [29]. In certain cases, our method is even more efficient in both training and rendering. Combined with the additional time required for video generation (as shown in Table 1), the overall time cost of our method remains acceptable.
4.6 Qualitative Comparison
We present the original video head frames, the corresponding stylized video frames generated by our method, and the 3D stylized head avatars synthesized using Gaussian blendshapes. The qualitative comparison is shown in Figure 5. The examples are selected from the “bala” dataset in the “Pixar” style and the “wojtek_1” dataset in the “Arcane” style.
In the stylized video, the “bala” data exhibits artifacts along the side edge of the head. We attribute this to the latent space distribution learned by StyleGAN, which tends to produce striped artifacts when the viewing angle falls outside the distribution covered by the training data. Notably, these artifacts are not present in the corresponding 3D stylized head avatars rendered by our method. Furthermore, the 3D stylized head avatars successfully preserve fine details from the stylized videos, such as the mole near the eye in the “wojtek_1” dataset. However, since the 3D avatar synthesis mainly focuses on the facial region, the neck area is typically blurred, as observed in both cases. This blurring leads to the lower quantitative performance, since the neck region is included in the evaluation.
The qualitative comparison with 3DGB [29] is presented in Figures 6. Our method effectively captures and preserves high-frequency details in the stylized videos. Compared to the state-of-the-art method, ToonifyGB can synthesize 3D stylized head avatars with comparable quality and detail.
4.7 Visualization
To better demonstrate the visual quality of our generated videos, we present several examples in Figure 3, and select two representative videos for detailed comparison in Figure 4. Specifically, we show real head frames from the “bala” and “wojtek_1” videos, as well as the corresponding heads of generated videos in the “Arcane” style.
The results demonstrate that key facial features, such as the beard, mouth shape, and even small details like the black mole above the eye in the lower right image, are well preserved after the stylization process. These details highlight the excellent performance of our method in terms of detail preservation and identity consistency.
@@ 4.7 Visualization
4.8 User Study
To more effectively evaluate the 3D stylized head avatars synthesized by our method, we conduct a user preference study. We collect 180 votes for both the “Arcane” and “Pixar” styles, respectively, with users rating their preferences using Likert scales [27] across three evaluation criteria: Style Consistency, Identity Preservation, and Overall Quality. Each criterion is assessed using a five-point scale: 1 for Very Dissatisfied, 2 for Dissatisfied, 3 for Neutral, 4 for Satisfied, and 5 for Very Satisfied.
Specifically, Style Consistency evaluates how well the stylized output aligns with the defining characteristics of the style; Identity Preservation measures whether the avatar retains the unique features and identity of the original character after stylization; and Overall Quality provides a comprehensive assessment of the visual appeal and overall quality of the synthesized avatar.
The results are presented in Table 5, indicating that most users show great satisfaction with the “Arcane” style, particularly in terms of Style Consistency (51.7% of users are very satisfied) and Overall Quality (41.7% of users are very satisfied). Although the rating for Identity Preservation is slightly lower, the average score remains favorable.
In addition, the “Pixar” style is also favored by users, particularly in Style Consistency (49.4% of users are very satisfied). For Identity Preservation and Overall Quality, the majority of users (over 60%) indicate satisfaction with our 3D stylized head avatars.
Figure 8: Limitation: We present side-view renderings synthesized by 3DGB [29] and our method.
4.9 Ablation Study
4.9.1 Face Alignment and Cropping
@@
4.9.2 Source Videos for Driving Animation
Compared with the architecture of 3DGB [29] that synthesizes 3D photo-realistic head avatars, our framework includes an additional Stage 1 to generate the stylized video. To demonstrate the importance of the generated stylized video in driving the animation, we compare the results of using the original input video (real face) versus our generated stylized video as the driving source, as shown in Figure 7.
@@ 4.9.2 Source Videos for Driving Animation
5 Limitation
Our method struggles to render side views of 3D stylized head avatars when the training data (i.e., input video) lacks side-view representations of the real head. As shown in Figure 8, we present side-view renderings synthesized by both 3DGB [29] and our method, and this limitation is also observed in the state-of-the-art methods. In fact, existing NeRF-based and Gaussian-based methods have yet to effectively address this issue. Rendering novel views from single-view training data remains an open problem for future research. Two directions to address this limitation include employing 2D GANs to synthesize videos with side views as additional training data, and enhancing the generalization ability of our model.
6 Conclusion
We propose a novel two-stage framework, named ToonifyGB, which utilizes Gaussian blendshapes to synthesize head animations in diverse styles from monocular videos. In Stage 1, the proposed method adopts an improved StyleGAN-based model to generate stylized videos without requiring face alignment or cropping as preprocessing. This results in more temporally stable outputs, providing a reliable foundation for high-quality 3D head avatar animation synthesis. Stage 2 focuses on constructing 3D stylized head avatars using Gaussian blendshapes, enabling fine-grained expression modeling and satisfactory animation. Our method supports real-time generation of stylized avatar animations in popular styles such as “Arcane” and “Pixar”.
For future work, we plan to integrate motion capture technologies to enable real-time expression control of 3D stylized avatars. Specifically, we aim to explore more efficient approaches for obtaining real-time expression parameters, bypassing the complexity of traditional PCA inversion. This direction is expected to further broaden the applicability of ToonifyGB in virtual character interaction and personalized avatar generation.
Acknowledgments
This work was supported in part by the National Science and Technology Council (NSTC), Taiwan, under Grants NSTC 114-2221-E-002-001- and NSTC 114-2420-H-002-010-.